# latexOverwriting file

### v4

@@ 2.1 The Algorithm: Gaussian Case
Here $\widehat{\mathit{C}}_{j+1}$ denotes the covariance of $\widehat{v}_{j+1}$. A simple calculation reveals that, if $v_{0}\sim\mathcal{N}(m_{0},\mathit{C}_{0})$, then in fact $\widehat{v}_{j+1}\sim\mathcal{N}(\widehat{m}_{j+1},\widehat{\mathit{C}}_{j+1})$ and $v_{j+1}\sim\mathcal{N}(m_{j+1},\mathit{C}_{j+1})$, where the means and covariances are given by the Kalman filter (1.4.2). Note that (2.1) constitutes a mean-field stochastic dynamical system because equation (2.1c) requires knowledge of $\widehat{\mathit{C}}_{j+1}$, which depends on the law $\mu_{j}$ of $v_{j}$. The law of this mean-field stochastic dynamical system is thus equal to the law of the Kalman filter. The analysis in [44, 46] is concerned with studying particle approximations of this Gaussian mean-field stochastic dynamical system, and convergence to the mean-field limit in the limit of infinite particles. In contrast, our work concerns the study of mean-field stochastic dynamical systems, but goes beyond the Gaussian setting. To this end, the next subsection defines the mean-field ensemble Kalman filter in the general, non-Gaussian, setting.
@@ 2.2 The Algorithm: General Case
The ensemble Kalman filter may be derived as a particle approximation of various mean field limits [8]. The specific mean field ensemble Kalman filter that we study in this paper reads
@@
where $\widehat{\pi}^{K}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and independence of the noise terms (1.3) is still assumed to hold. See Subsection 1.3 for the definition of the covariance matrices that appear in (2.2c). Let us denote by $\mu^{K}_{j}$ the law of $u_{j}$. In order to rewrite the evolution of $\mu_{j}^{K}$ in terms of maps on probability measures, let us introduce for a given $y^{\dagger}_{j+1}$ the map $\mathsf{T}_{j}\colon\mathcal{P}(^{d}\times^{K})\to\mathcal{P}(^{d})$ defined by
@@
Here the subscript ${}_{\sharp}$ denotes the pushforward and, for any given $\pi\in\mathcal{P}(^{d}\times^{K})$ and $z\in^{K}$, the map $\mathscr{T}$ is affine in its first two arguments and given by
@@
As demonstrated in [8], with this notation the evolution of the probability measure $\mu_{j}^{K}$ may be written in compact form as
@@
    $B_{y}:=\left\{Y^{\dagger}\in^{KJ}:\max_{j\in}\lvert y^{\dagger}_{j}\rvert\leqslant\kappa_{y}\right\}.$
@@
    The function $h$ satisfies $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\in B_{h}:=B_{L^{\infty}}(0,\kappa_{h})$ for some $\kappa_{h}>0$.
@@ 2.3 Stability Theorem: Ensemble Kalman Filter
Roughly speaking, our main result states that if the true filtering distributions $(\mu_{j})_{j\in}$ are close to Gaussian after appropriate lifting to the state/data space then the distribution $\mu^{K}_{j}$ given by the mean field ensemble Kalman filter (2.5) is close to the true filtering distribution $\mu_{j}$ given by (1.9) for all $j\in\llbracket 0,J\rrbracket$. Recall that $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ denotes the $C^{0,1}$ semi-norm, namely the Lipschitz constant.
@@
Assume that the probability measures $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ are obtained respectively from the dynamical systems (1.9) and (2.5), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,
@@
$d_{g}(\mu^{K}_{J},\mu_{J})\leqslant C\max_{j\in}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
@@
The stability theorem shows that the error made by the ensemble Kalman filter is controlled by the difference between the true filter, lifted to the joint space of state and observations, and its Gaussian projection. The constant $C$ in the stability estimate is uniform across realizations of the data from $B_{y}.$ Indeed, since $h$ is assumed bounded, the bound $\kappa_{y}$ on the data will hold with probability exponentially close to $1$, [^1] With respect to the noise realization leading to the data. for $\kappa_{y}\gg 1$, when there is no model mis-specification. Relaxing the assumptions of bounded $\Psi$ and $h$, for example to the setting of linear plus bounded functions, is technically challenging and will be left for future work. Likewise relaxing positivity assumptions on the noise leads to substantial technical obstacles, deferred for future work. And finally, relaxing the assumption that $J$ is finite will require further structural assumptions on the long-time behaviour of the filter, and is also deferred for future work.
@@
    For any probability measure $\mu$, the first moments of the probability measures $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their second moments are bounded both from above and from below. The constants in these bounds depend only on the parameters $\kappa_{\Psi}$, $\kappa_{h}$, $\mathit{\Sigma}$ and $\mathit{\Gamma}$. See Lemmas 15 and 16.
@@
    For any Gaussian measure $\mu\in\mathcal{G}(^{d}\times^{K})$, it holds that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma 19 and also [8].
@@
$\varepsilon:=\max_{j\in}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$ (2.6)
Assume throughout the following that $j\in.$ The main idea of the proof results from the following use of the triangle inequality
@@
We have used the fact that the second argument of the second term on the right-hand side indeed coincides with the first argument of the third term because $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by item 2 (Lemma 19.) By item 1 (Lemma 16), there is a constant $R\geqslant 1$, depending on the covariance matrices $\mathit{\Sigma}$, $\mathit{\Gamma}$ and the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, such that $\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}_{R}(^{d}\times^{K})$. By items 3, 4 and 6 (Lemmas 24, 22 and 21), the first term in (2.7b) satisfies
@@ 2.4 Approximation Theorem: Ensemble Kalman Filter
Let $(\mathit{\Sigma},\mathit{\Gamma})$ satisfy Assumption (H4). Suppose that $\Psi_{0}\colon^{d}\to^{d}$ and $h_{0}\colon^{d}\to^{K}$ are functions taking constant values and denote by $B_{\Psi_{0},h_{0}}(r)$ the set of all functions $(\Psi,h)$ satisfying $\Psi\in B_{L^{\infty}}(\Psi_{0},r)$, $h\in B_{L^{\infty}}(h_{0},r)$ and Assumptions (H2) and (H3). Assume also that $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$ and denote by $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ the probability measures obtained respectively from the dynamical systems (1.9) and (2.5), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,
@@ 3 Generalization: Gaussian Projected Filter
We generalize the main result to the Gaussian projected filter defined in [8]. This algorithm may be viewed as a mean field version of the unscented Kalman filter [37]. Using a similar approach to that adopted in the previous section, we prove a similar stability theorem and accuracy corollary. The Gaussian projected filter is defined by the iteration
@@
Assume that the probability measures $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ are obtained respectively from the dynamical systems (1.9) and (3.1), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{G}$. That is,
@@
$d_{g}(\mu^{G}_{J},\mu_{J})\leqslant C\max_{j\in}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
The proof of this error estimate relies on the auxiliary results Lemmas 16, 21, 22 and 23 already discussed, as well as the following two additional results.
@@
  • The map $\mathsf{B}_{j}$ is locally Lipschitz for the metric $d_{g}$ over Gaussians, in the sense that for any $R\geqslant 1$ and any $j\in$, this map is Lipschitz continuous over $\mathcal{G}_{R}(^{d}\times^{K})$. The associated Lipschitz constant is denoted by $L_{\mathsf{B}}=L_{\mathsf{B}}(R,\kappa_{y})$. See Lemma 29.
@@
It follows from Lemmas 16, 21, 22, 26 and 29 that the composition of maps $\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}$ is globally Lipschitz continuous on $\mathcal{P}(^{d})$ with a constant $\ell=\ell(R,\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$, where $R=R(\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$ is a positive constant such that $\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}_{R}(^{d}\times^{K})$. Therefore, the first term on the right hand side may be bounded by
@@
Let $(\mathit{\Sigma},\mathit{\Gamma})$ satisfy Assumption (H4). Suppose that $\Psi_{0}\colon^{d}\to^{d}$ and $h_{0}\colon^{d}\to^{K}$ are functions taking constant values and denote by $B_{\Psi_{0},h_{0}}(r)$ the set of all functions $(\Psi,h)$ satisfying $\Psi\in B_{L^{\infty}}(\Psi_{0},r)$, $h\in B_{L^{\infty}}(h_{0},r)$ and Assumptions (H2) and (H3). Assume that the probability measures $(\mu_{j})_{j\in}$ and $(\mu^{K}_{j})_{j\in}$ are obtained respectively from the dynamical systems (1.9) and (3.1), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,

### v6

@@ 2.1 The Algorithm: Gaussian Case
Here $\widehat{\mathit{C}}_{j+1}$ denotes the covariance of $\widehat{v}_{j+1}$. A simple calculation reveals that, if $v_{0}\sim\mathcal{N}(m_{0},\mathit{C}_{0})$, then in fact $\widehat{v}_{j+1}\sim\mathcal{N}(\widehat{m}_{j+1},\widehat{\mathit{C}}_{j+1})$ and $v_{j+1}\sim\mathcal{N}(m_{j+1},\mathit{C}_{j+1})$, where the means and covariances are given by the Kalman filter (1.10). Note that (2.1) constitutes a mean-field stochastic dynamical system because equation (2.1c) requires knowledge of $\widehat{\mathit{C}}_{j+1}$, which depends on the law $\mu_{j}$ of $v_{j}$. The law of this mean-field stochastic dynamical system is thus equal to the law of the Kalman filter. The analysis in [le2009large, mandel2011convergence] is concerned with studying particle approximations of this Gaussian mean-field stochastic dynamical system, and convergence to the mean-field limit in the limit of infinite particles. In contrast, our work concerns the study of mean-field stochastic dynamical systems, but goes beyond the Gaussian setting. To this end, the next subsection defines the mean-field ensemble Kalman filter in the general, non-Gaussian, setting.
@@ 2.2 The Algorithm: General Case
The ensemble Kalman filter may be derived as a particle approximation of various mean field limits [2022arXiv220911371C]. The specific mean field ensemble Kalman filter that we study in this paper reads
@@
where $\widehat{\pi}^{K}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and independence of the noise terms (1.3) is still assumed to hold. The covariance matrices $\mathit{\Gamma},\mathit{\Sigma}$ are still assumed to be positive definite, so $\mathcal{C}^{yy}(\widehat{\pi}_{j+1}^{K})\succ 0$ and the algorithm is well-defined. See Subsection 1.3 for the definition of the covariance matrices that appear in (2.2c).
Remark 3.
The mean-field EnKF algorithm may be derived as the best linear unbiased estimator (BLUE) of the predicted state, given the data; see [2022arXiv220911371C].
Let us denote by $\mu^{K}_{j}$ the law of $u_{j}$. In order to rewrite the evolution of $\mu_{j}^{K}$ in terms of maps on probability measures, let us introduce
$\mathcal{P}^{2}_{\succ 0}\left({}^{d}\times^{K}\right):=\Bigl{\{}\pi\in\mathcal{P}^{2}\left({}^{d}\times^{K}\right):\mathcal{C}^{yy}(\pi)\succ 0\Bigr{\}}.$
Then, for a given $y^{\dagger}_{j+1}$, we denote by $\mathsf{T}_{j}\colon\mathcal{P}^{2}_{\succ 0}(^{d}\times^{K})\to\mathcal{P}^{2}(^{d})$ the map defined by
@@
Here the subscript denotes the pushforward and, for any given $\pi\in\mathcal{P}(^{d}\times^{K})$ and $z\in^{K}$, the map $\mathscr{T}$ is affine in its first two arguments and given by
@@
As demonstrated in [2022arXiv220911371C], with this notation the evolution of the probability measure $\mu_{j}^{K}$ may be written in compact form as
@@
    $B_{y}:=\left\{Y^{\dagger}\in^{KJ}:\max_{j\in\llbracket 0,J\rrbracket}\lvert y^{\dagger}_{j}\rvert\leqslant\kappa_{y}\right\}.$
@@
    The function $h$ is continuous and satisfies $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\in B_{h}:=B_{L^{\infty}}(0,\kappa_{h})$ for some $\kappa_{h}>0$.
@@ 2.3 Stability Theorem: Ensemble Kalman Filter
Roughly speaking, our main result states that if the true filtering distributions $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ are close to Gaussian after appropriate lifting to the state/data space then the distribution $\mu^{K}_{j}$ given by the mean field ensemble Kalman filter (2.5) is close to the true filtering distribution $\mu_{j}$ given by (1.9) for all $j\in\llbracket 0,J\rrbracket$. Recall that $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ denotes the $C^{0,1}$ semi-norm, namely the Lipschitz constant.
@@
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{K}_{j})_{j\in\llbracket 1,J\rrbracket}$ are obtained respectively from the dynamical systems (1.9) and (2.5), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,
@@
$d_{g}(\mu^{K}_{J},\mu_{J})\leqslant C\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
@@
The constant $C$ in the stability estimate is uniform across realizations of the data from $B_{y}.$ Indeed, since $h$ is assumed bounded, the bound $\kappa_{y}$ on the data will hold with probability exponentially close to $1$, [^1] With respect to the noise realization leading to the data. for $\kappa_{y}\gg 1$, when there is no model mis-specification. Relaxing the assumptions of bounded $\Psi$ and $h$, for example to the setting of linear plus bounded functions, is technically challenging and will be left for future work. Likewise relaxing positivity assumptions on the noise leads to substantial technical obstacles, deferred for future work. And finally, relaxing the assumption that $J$ is finite will require further structural assumptions on the long-time behaviour of the filter, and is also deferred for future work.
Remark 6.
The stability theorem shows that the error made by the ensemble Kalman filter is controlled by the difference between the true filter, lifted to the joint space of state and observations, and its Gaussian projection. To interpret the result it is thus important to have intuition about what it means for a measure to be close to its Gaussian projection in the $d_{g}$ metric. To this end note that, by Remark 2, it is necessary that the means and covariances of the measure and its Gaussian projection are close, for the measures to be close in the $d_{g}$ metric. This is automatically satisfied for the difference between a measure and its Gaussian projection, by construction. Since closeness in $d_{g}$ requires the expectations of all functions growing no faster than quadratic to be close, a useful rule of thumb for practitioners is that the quantity $d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j})$ is small when matching first and second moments allows control of all quadratically bounded functions. This will happen, for example, when $\Psi$ and $h$ are small nonlinear perturbations of affine functions; and hence it will happen when $\Psi$ and $h$ are small nonlinear perturbations of either linear functions or constant functions. For technical reasons (see previous remark) this paper excludes the case of linearly growing functions, but we do study small perturbations of constant functions in Corollary 7. There is also reason to expect the filtering distribution to be close to Gaussian when the data volume is high, or noise is small, and some version of observability applies; then central limit theorems for inverse problems, such as the Bernstein von Mises theorem, may in the future be developed to quantify this assertion [van2000asymptotic]; for work pointing in this direction, but in the infinite dimensional case where Bernstein von Mises type results are harder to establish, see [nickl2023posterior].
@@
    For any probability measure $\mu$, the first moments of the probability measures $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their second moments are bounded both from above and from below. The constants in these bounds depend only on the parameters $\kappa_{\Psi}$, $\kappa_{h}$, $\mathit{\Sigma}$ and $\mathit{\Gamma}$. See Lemmas 17 and 18.
@@
    For any Gaussian measure $\mu\in\mathcal{G}(^{d}\times^{K})$, it holds that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma 21 and also [2022arXiv220911371C].
@@
$\varepsilon:=\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$ (2.6)
Assume throughout the following that $j\in\llbracket 0,J-1\rrbracket.$ The main idea of the proof results from the following use of the triangle inequality
@@
We have used the fact that the second argument of the second term on the right-hand side indeed coincides with the first argument of the third term because $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by item 2 (Lemma 21.) By item 1 (Lemma 18), there is a constant $R\geqslant 1$, depending on the covariance matrices $\mathit{\Sigma}$, $\mathit{\Gamma}$ and the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, such that $\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}_{R}(^{d}\times^{K})$. By items 3, 4 and 6 (Lemmas 26, 24 and 23), the first term in (2.7b) satisfies
@@ 2.4 Approximation Theorem: Ensemble Kalman Filter
Let $(\mathit{\Sigma},\mathit{\Gamma})$ satisfy Assumption (H4). Suppose that $\Psi_{0}\colon^{d}\to^{d}$ and $h_{0}\colon^{d}\to^{K}$ are functions taking constant values and denote by $B_{\Psi_{0},h_{0}}(r)$ the set of all functions $(\Psi,h)$ satisfying $\Psi\in B_{L^{\infty}}(\Psi_{0},r)$, $h\in B_{L^{\infty}}(h_{0},r)$ and Assumptions (H2) and (H3). Assume also that $|h|_{C^{0,1}}\leqslant\ell_{h}<\infty$ and denote by $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{K}_{j})_{j\in\llbracket 1,J\rrbracket}$ the probability measures obtained respectively from the dynamical systems (1.9) and (2.5), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,
@@ 3 Generalization: Gaussian Projected Filter
We generalize the main result to the Gaussian projected filter defined in [2022arXiv220911371C]. This algorithm may be viewed as a mean field version of the unscented Kalman filter [julier1997new]. Using a similar approach to that adopted in the previous section, we prove a similar stability theorem and accuracy corollary. The Gaussian projected filter is defined by the iteration
@@
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{K}_{j})_{j\in\llbracket 1,J\rrbracket}$ are obtained respectively from the dynamical systems (1.9) and (3.1), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{G}$. That is,
@@
$d_{g}(\mu^{G}_{J},\mu_{J})\leqslant C\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}).$
The proof of this error estimate relies on the auxiliary results Lemmas 18, 23, 24 and 25 already discussed, as well as the following two additional results.
@@
  • The map $\mathsf{B}_{j}$ is locally Lipschitz for the metric $d_{g}$ over Gaussians, in the sense that for any $R\geqslant 1$ and any $j\in\llbracket 0,J-1\rrbracket$, this map is Lipschitz continuous over $\mathcal{G}_{R}(^{d}\times^{K})$. The associated Lipschitz constant is denoted by $L_{\mathsf{B}}=L_{\mathsf{B}}(R,\kappa_{y})$. See Lemma 31.
@@
It follows from Lemmas 18, 23, 24, 28 and 31 that the composition of maps $\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}$ is globally Lipschitz continuous on $\mathcal{P}(^{d})$ with a constant $\ell=\ell(R,\kappa_{y},\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$, where $R=R(\kappa_{\Psi},\kappa_{h},\mathit{\Sigma},\mathit{\Gamma})$ is a positive constant such that $\operatorname{Im}(\mathsf{Q}\mathsf{P})\subset\mathcal{P}_{R}(^{d}\times^{K})$. Therefore, the first term on the right hand side may be bounded by
@@
Let $(\mathit{\Sigma},\mathit{\Gamma})$ satisfy Assumption (H4). Suppose that $\Psi_{0}\colon^{d}\to^{d}$ and $h_{0}\colon^{d}\to^{K}$ are functions taking constant values and denote by $B_{\Psi_{0},h_{0}}(r)$ the set of all functions $(\Psi,h)$ satisfying $\Psi\in B_{L^{\infty}}(\Psi_{0},r)$, $h\in B_{L^{\infty}}(h_{0},r)$ and Assumptions (H2) and (H3). Assume that the probability measures $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{K}_{j})_{j\in\llbracket 1,J\rrbracket}$ are obtained respectively from the dynamical systems (1.9) and (3.1), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{K}\in~{}\mathcal{G}(^{d})$. That is,